write the equation for two times the sum of a number and three is equal to fifteen.
I wrote 2(n+3)=15 but do not think it is correct because when I checked the answer into the equation my answer did not equal fifteen.

Answer by edjones(8007)

2(n+3)=15
2n+6=15
2n=9
n=4.5
.
Check:
2(4.5+3)=2*7.5=15
.
Ed